OpenStudy (anonymous):

5(2)^(3x)-4=13 solve for x

OpenStudy (anonymous):

?????

OpenStudy (cwrw238):

5(2)^(3x)-4=13 5(2)^(3x) = 17 2^(3x) = 17/5 take logarithms of both sides: 3x * ln2 = ln 3.4 con u continue from here?

OpenStudy (anonymous):

no

OpenStudy (cwrw238):

ok divide both sides by ln 2 3x = ln 3.4 / ln2 x = ln3.4 / (3 ln 2 )

OpenStudy (anonymous):

thats the answer?

OpenStudy (cwrw238):

you need to compute ln 3.4 / (3 ln 2) with your calculator

OpenStudy (anonymous):

in radian or degree mode?

OpenStudy (cwrw238):

whenever you have the variable in the index take logs

OpenStudy (cwrw238):

it doesn't matter unless x is an angle

OpenStudy (anonymous):

ok